What is the equation of the line passing through the origin and parallel to the line passing through the points (2,8) and (4,12)?
Answer by lwsshak3(11628)

What is the equation of the line passing through the origin and parallel to the line passing through the points (2,8) and (4,12)?
***
equation of a line: y=mx+b, m=slope, b=y-intercept
slope of line passing through the points (2,8) and (4,12)?
slope=∆y/∆x=(4-2)/(12-8)=2/4=1/2
slope of parallel line=1/2
y-intercept of parallel line=0
equation of parallel line: y=x/2